Baddie156:

Ohm’s law is I = V ÷ R. What’s the voltage of the battery described? A circuit has one 10.0-ohm resistor. The current flowing through the circuit is 0.15 amps. So, the battery in the circuit is a -volt battery.

Baddie156:

plz help

AZ:

You have the formula and you have the values Ohm's law is V = IR Voltage = Current * Resistance Current is 0.15 A Resistance is 10.0 ohms Voltage is what we're solving for So if we put in the numbers into the equation, what do we get? Voltage = 0.15 * 10.0 Voltage = ??

Baddie156:

According to Ohm’s law, I = V ÷ R. Rearrange the equation: V = I × R. Then plug in the values from the problem statement: V = 0.15 × 10.0 V = 1.5. The battery in the circuit is a 1.5-volt battery.

AZ:

That's correct!

Baddie156:

omg thx

AZ:

Of course!